OpenStudy (anonymous):

Can you help me integrate function (sinx cosx)/((sinx)^3+(cosx)^3)?

OpenStudy (anonymous):

Put u = tan(x/2)

OpenStudy (anonymous):

Then du = 1/2sec^2(x/2)dx

OpenStudy (anonymous):

Now use:\[\sin x = 2u/(u^2+1) ,......\cos u=(1-u^2)/(u^2+1)\]

OpenStudy (anonymous):

Thank you:)

OpenStudy (anonymous):

Welcome!